# thanks: The first two authors contributed equally

### v1

@@ Abstract
The memory effect, in the context of gravitational-waves (GWs), manifests itself in the permanent relative displacement of test masses when they encounter the GWs. A number of works have explored the possibility of detecting the memory when the source and detector are separated by large distances. A special type of memory, arising from BMS symmetries, called “black-hole memory”, has been recently proposed. The black hole memory only manifests itself in the vicinity of its event horizon. Therefore, formally observing it requires placing a GW detector at the horizon of the BH, which prima-facie seems unfeasible. In this work, we describe a toy model that suggests a possible way the black hole memory may be observed, without requiring a human-made detector near the event horizon. The model considers a binary black hole (BBH), emanating GWs observable at cosmological distances, as a proxy for an idealized detector in the vicinity of a supermassive Schwarzschild black hole that is endowed with a supertranslation hair by sending a shock-wave to it. This sudden change affects the geometry near the horizon of the supertranslated black hole and it induces a change in the inspiraling orbital separation (and hence, orbital frequency) of the binary, which in turn imprints itself on the GWs. Using basic GW data analysis tools, we demonstrate that the black hole memory should be observable by a LISA-like space-based detector.
@@ I Introduction
The LIGO-Virgo (LV) detector network Aasi et al. (2015); Acernese et al. (2015) has detected $\sim 90$ GW events produced by compact binary coalescences (CBCs), across three observing runs (O1, O2, O3) Abbott et al. (2023). The majority of these are binary black hole (BBH) mergers, and have therefore enabled some of the most unique tests of general relativity (GR) in the strong field regime Abbott et al. (2021). These include a model-agnostic test of GR which subtracts out the best matched GR template from the data known to contain the GW signal, and examines if the residual is consistent with noise Abbott et al. (2016); Abbott et al. (2019a); Abbott et al. (2020); an inspiral-merger-ringdown consistency test that checks for consistency between the low and high frequency portions of the GW waveform Ghosh et al. (2016, 2018); a test that looks for deviations in the Post-Newtonian (PN) parameters governing the inspiral evolution of the CBC Blanchet and Sathyaprakash (1994, 1995); Arun et al. (2006a, b); Yunes and Pretorius (2009); Mishra et al. (2010); Li et al. (2012a, b); and propagation tests that compare the speed of GWs with the speed of light Abbott et al. (2019b), as well as signatures of velocity dispersion due to a finite graviton mass Will (1998).
@@
This remarkable feature of spacetime is a genuine test of general relativity. The linear memory appears due to the gravitational radiation generated by binaries executing unbounded (hyperbolic) trajectories and is generally expected to be vanishing for binaries with bound orbits.  [^1] Although for bound orbits some contribution from linear, nonheriditary DC terms appears at 5PN and higher orders, it is not clear that these effects will lead to a permanent memory Favata (2009). Nevertheless, this effect is very weak and has a very low detection prospect. However, there is another type of memory that originates at the 2.5PN order in the radiative multipoles for quasi-circular and elliptical binary inspirals but contributes at the 0PN order in the waveform Favata (2009). This is in contrast to the hyperbolic, parabolic and radial orbits where this effect occurs at 2.5PN order relative to the leading order (Newtonian) term in the waveform Favata (2011). This is known as non-linear memory or Christodoulou memory Christodoulou (1991); Blanchet and Damour (1992); Thorne (1992). This memory has a simple interpretation: it is the “wave of a wave”, i.e, it arises due to GWs sourced by GWs. This memory signal is expected to be detected by future space-based detectors like LISA (see, e.g., Johnson et al. (2019)).
Recently, an interesting relation between the memory effect and asymptotic symmetries of spacetimes has been established. The asymptotic symmetries are the symmetries that preserve the asymptotic structure of spacetimes. This was first discovered by Bondi-van-der Burg-Metzner and Sachs (BMS) and described in their pioneering works Bondi et al. (1962); Sachs (1962). The group of symmetries recovered near the infinity of an asymptotically flat spacetime is termed as BMS group. In its primitive form, the BMS group contained an infinite dimensional abelian subgroup whose generators are called supertranslations. These are angle dependent translations that generalize the ordinary rigid translations of flat spacetime. Later, the asymptotic group has been extended to accommodate superrotations and can be viewed as conformal symmetries of the celestial sphere at null infinity ($\mathcal{I}^{+}$) Barnich and Troessaert (2010); Barnich and Troessaert (2012).
@@
Such conservation laws should also be true in the presence of asymptotically flat black holes. However, in this case, the charge conservation requires the black holes should also carry supertranslation (or superrotations) charges or some (soft) ‘hairs’ Hawking et al. (2017); Strominger (2018). In the presence of black holes, the future null infinity ${\mathcal{I}}^{+}$ fails to act as a Cauchy surface. One needs to add the future horizon ${\mathcal{H}}^{+}$ also, and ${\mathcal{I}}^{+}\cup{\mathcal{H}}^{+}$ acts as a Cauchy surface. Therefore, it is desirable to recover the BMS symmetries near the horizon of a black hole in spacetime as well.
BMS symmetries have been recovered near the horizon of a black hole by several theoretical approaches. BMS-like symmetries can be discovered at the null boundaries including black hole horizons by analyzing diffeomorphisms that preserve certain universal structures of these surfaces Chandrasekaran et al. (2018); Ashtekar et al. (2018). In another approach, near-horizon asymptotic symmetries are recovered by preserving the near-horizon asymptotic structures Koga (2001); Donnay et al. (2016a, b) [^2] BMS symmetries have also been recovered as soldering symmetries of thin shells. Readers may look at Blau and O’Loughlin (2016); Bhattacharjee and Bhattacharyya (2018) and the references therein.
@@
Clearly, the existence of such a near-horizon memory or ‘hair’ can only be detected by an idealized observer hovering around the horizon. Can such a memory be recognised from the far region? Particularly, would it be possible to think of a scenario in which the near-horizon memory can show up in future GW detectors? In this article, we propose a toy model that can provide plausible answers to these questions.
@@ II Implantation of memory and its detection
It is apparent that to detect the near-horizon supertranslation memory, one needs to place the detectors close to the horizon. However, detection from a far region would not be possible until the detectors can have some mechanism to communicate with the remote region. To this end, we consider a simplified model and indicate a possible way of obtaining imprints of such near-horizon memory in GWs detectable by the upcoming space-based detectors like LISA.
II.1 The Black Hole Memory
The detection setup uses a toy model in which an eternal Schwarzschild black hole is endowed with a linearised supertranslation hair by throwing some asymmetric shockwave into it as proposed in Hawking et al. (2017); Strominger (2018). This process implants a linearized supertranslation hair to a bald Schwarzschild black hole:
@@ II.1 The Black Hole Memory
This change in the transverse part of the metric components gives rise to memory for an observer hovering around the horizon. As the physical degrees of freedom are encoded in the transverse part of the metric, this produces a genuine effect.
BMS-like symmetries can also be recovered near the horizon of a black hole that preserves the near-horizon asymptotic structure. To see this, one may expand the metric in Eq. (II.1) near the horizon by a suitably defined radial coordinate $\rho=r-r^{h}_{(f)}$ and relate the parameters of the near-horizon metric with the implanted supertranslation field Donnay et al. (2016b, 2018). To the leading order in $\rho$ this metric reads:
@@
The changes induced via horizon supertranslation (or superrotation) are termed as ‘black hole memory effect’ Donnay et al. (2018). The change induced by the shock-wave can be realized as the action of supertranslation and superroation on the near-horizon geometry. One can obtain the charges corresponding to these near-horizon symmetries. The supertranslation and superrotation charges depend upon the field $f$ or its derivatives, 333see Donnay et al. (2016b) for details. and therefore, if we detect $f$ by any mechanism, we detect the supertranslation or the black hole memory.
@@ II.2 A toy detection model
Figure 1: Schematic representation of inspiralling IMBBH source emitting GWs in the vicinity of the horizon of a bald SMBH (on the left). Once the supertranslation (ST) ‘hair’ is attached to the bald SMBH due to some astrophysical shock wave, the proper distance between the components of the IMBBH is changed. The magnitude of the change in separation (ST memory) denoted by $\Delta L$ is given by (11). As the IMBBH separation increases (decreases), the binary will start emitting GWs at a lower (higher) frequency in comparison to the initial configuration (on the left).
@@
Now, we propose a scenario (probe for detection) that might capture this supertranslation memory. Ideally, to detect such a memory one would need to place the detectors near the vicinity of the horizon. Suppose two intermediate-mass black holes (IMBHs) are in the vicinity of a bald supermassive black hole (SMBH) and they act like detectors. The IMBHs initially keep generating GWs which are detectable by LISA. Now suddenly the SMBH (almost instantaneously) gets endowed with a (linear) supertranslation hair by sending a shockwave into it as described above. This sudden effect would change the relative proper distance between the inspiraling IMBHs and would ultimately alter the inspiraling frequency and frequency evolution. Note that the shockwave acts only as an agent that implants the supertranslation hair on the bald SMBH. This is a convenient way to generate a BH state with a (asymptotic) supertranslation field.
@@
We consider the IMBHs to be separated only in the angular directions initially, i.e. they are at the same radial distance from the centre of the hole. After the hair is implanted, the separation between the black holes changes. This can be estimated using local coordinates. We will further assume the supertranslation field to be a function of $\theta$ only. This would reduce the algebra and at the same time offer sufficient insights to the desired investigation. The metric (cf. Eq. II.1) can now be cast in a more explicit form as:
$ds^{2}=$ $-{1\over 2M}\rho dv^{2}+2dvd\rho-{\rho\over M}\partial_{\theta}fd\theta dv$
$+$ $(4M^{2}d\theta^{2}+4M^{2}\sin\theta^{2}d\phi^{2}+4M\partial_{\theta}^{2}fd\theta^{2}).$ (9)
@@
Now, let $L_{0}$ be the initial proper distance (angular) between the test bodies (IMBHs). In that case, the separation would only be governed by the bald angular part of the metric. Therefore $L_{0}=4M^{2}\gamma_{AB}d\Theta^{A}d\Theta^{B}.$ After the implantation of the supertranslation memory, the changed distance would be governed by the metric given by Eq. (II.2). If $\Delta L$ is the shift in the separation then a straightforward computation yields:
$(\Delta L)_{\mathrm{exact}}=\sqrt{L_{0}^{2}-12Mq\cos(2\theta)(\delta\theta)^{2}}-L_{0},$ (10)
@@
where we have chosen $f(\theta)$ to be equal to the second Legendre Polynomial $f=qP_{2}(\cos\theta).$ $q$ is a parameter that encapsulates the strength of the supertranslation hair Sarkar et al. (2022); Compère and Long (2016). For the case under consideration, $q$ should be small as it is a part of a weak supertranslation field. Practically the initial separation $L_{0}$ should be much greater than the separation induced by the field $f$, so one gets:
$\Delta L\simeq-{6M\over L_{0}}q\cos(2\theta)(\delta\theta)^{2}.$ (11)
Here, $\delta\theta$ is the initial azimuthal separation between the objects. The shift in the separation of IMBHs is parametrized by $q$, which will induce a change in the frequency of their inspiraling motion.
II.3 Frequency modulation of the binary

### v2

@@ Abstract
The memory effect, in the context of gravitational-waves (GWs), manifests itself in the permanent relative displacement of test masses when they encounter the GWs. A number of works have explored the possibility of detecting the memory when the source and detector are separated by large distances. A special type of memory, arising from Bondi–Metzner–Sachs (BMS)-like symmetries, called “black-hole memory”, has been recently proposed. The black hole memory only manifests itself in the vicinity of its event horizon. Therefore, formally observing it requires placing a GW detector at the horizon of the BH, which prima-facie seems unfeasible. In this work, we describe toy models that suggest a possible way the black hole memory may be observed, without requiring a human-made detector near the event horizon. The models consider a binary black hole (BBH), emanating GWs observable at cosmological distances, as a proxy for an idealized detector in the vicinity of a supermassive black hole that is endowed with a supertranslation hair by sending a shock-wave to it. This sudden change affects the geometry near the horizon of the supertranslated black hole and it induces a change in the inspiraling orbital separation (and hence, orbital frequency) of the binary, which in turn imprints itself on the GWs. Using basic GW data analysis tools, we demonstrate that the black hole memory should be observable by a LISA-like space-based detector.
@@ I Introduction
The LIGO-Virgo (LV) detector network Aasi et al. (2015); Acernese et al. (2015) has detected $\sim 90$ GW events produced by compact binary coalescences (CBCs), across three observing runs (O1, O2, O3) Abbott et al. (2023). The majority of these are binary black hole (BBH) mergers, and have therefore enabled some of the most unique tests of general relativity (GR) in the strong field regime Abbott et al. (2021). These include a model-agnostic test of GR which subtracts out the best matched GR template from the data known to contain the GW signal, and examines if the residual is consistent with noise Abbott et al. (2016, 2019a, 2020); an inspiral-merger-ringdown consistency test that checks for consistency between the low and high frequency portions of the GW waveform Ghosh et al. (2016, 2018); a test that looks for deviations in the Post-Newtonian (PN) parameters governing the inspiral evolution of the CBC Blanchet and Sathyaprakash (1994, 1995); Arun et al. (2006a, b); Yunes and Pretorius (2009); Mishra et al. (2010); Li et al. (2012a, b); and propagation tests that compare the speed of GWs with the speed of light Abbott et al. (2019b), as well as signatures of velocity dispersion due to a finite graviton mass Will (1998).
@@
This remarkable feature of spacetime is a genuine test of general relativity. The linear memory appears due to the gravitational radiation generated by binaries executing unbounded (hyperbolic) trajectories and is generally expected to be vanishing for binaries with bound orbits.  [^1] Although for bound orbits some contribution from linear, nonheriditary DC terms appears at 5PN and higher orders, it is not clear that these effects will lead to a permanent memory Favata (2009a). Nevertheless, this effect is very weak and has a very low detection prospect. However, there is another type of memory that originates at the 2.5PN order in the radiative multipoles for quasi-circular and elliptical binary inspirals but contributes at the 0PN order in the waveform Favata (2009a). This is in contrast to the hyperbolic, parabolic and radial orbits where this effect occurs at 2.5PN order relative to the leading order (Newtonian) term in the waveform Favata (2011). This is known as non-linear memory or Christodoulou memory Christodoulou (1991); Blanchet and Damour (1992); Thorne (1992). This memory has a simple interpretation: it is the “wave of a wave”, i.e, it arises due to GWs sourced by GWs. The detection prospects of nonlinear memory by ground-based and space-based detectors has been extensively studied Favata (2009b, a, c, 2010, 2011); Lasky et al. (2016); Grant and Nichols (2023); Ghosh et al. (2023). This memory signal is expected to be detected by future space-based detectors like Laser Interferometer Space Antenna (LISA) (see, e.g., Favata (2009c); Johnson et al. (2019); Ghosh et al. (2023)).
Recently, an interesting relation between the memory effect and asymptotic symmetries of spacetimes has been established. The asymptotic symmetries are the symmetries that preserve the asymptotic structure of spacetimes. This was first discovered by Bondi-van-der Burg-Metzner and Sachs (BMS) and described in their pioneering works Bondi et al. (1962); Sachs (1962). The group of symmetries recovered near the infinity of an asymptotically flat spacetime is termed as BMS group. In its primitive form, the BMS group contained an infinite dimensional abelian subgroup whose generators are called supertranslations. These are angle dependent translations that generalize the ordinary rigid translations of flat spacetime. Later, the asymptotic group has been extended to accommodate superrotations and can be viewed as conformal symmetries of the celestial sphere at null infinity ($\mathcal{I}^{+}$) Barnich and Troessaert (2010, 2012).
@@
Such conservation laws should also be true in the presence of asymptotically flat black holes. However, in this case, the charge conservation requires the black holes should also carry supertranslation (or superrotations) charges or some (soft) ‘hairs’ Hawking et al. (2017); Strominger (2018). In the presence of black holes, the future null infinity $\mathcal{I}$$^{+}$ fails to act as a Cauchy surface. One needs to add the future horizon $\mathcal{H}$$^{+}$ also, and $\mathcal{I}$$^{+}\cup{\mathcal{H}}^{+}$ acts as a Cauchy surface. Therefore, it is desirable to recover the BMS-like symmetries near the horizon of a black hole in spacetime as well.
BMS-like symmetries have been recovered near the horizon of a black hole by several theoretical approaches. BMS-like symmetries can be discovered at the null boundaries including black hole horizons by analyzing diffeomorphisms that preserve certain universal structures of these surfaces Chandrasekaran et al. (2018); Ashtekar et al. (2018). In another approach, near-horizon asymptotic symmetries are recovered by preserving the near-horizon asymptotic structures Koga (2001); Donnay et al. (2016a, b) [^2] BMS-like symmetries have also been recovered as soldering symmetries of thin shells. Readers may look at Blau and O’Loughlin (2016); Bhattacharjee and Bhattacharyya (2018) and the references therein.
@@
Clearly, the existence of such a near-horizon memory can only be detected by an idealized observer hovering around the horizon. Can such a memory be recognised from the far region? Particularly, would it be possible to think of a scenario in which the near-horizon memory can show up in future GW detectors? In this article, we propose a toy model that can provide plausible answers to these questions.
@@ II Implantation of memory and its detection
We consider two scenarios to detect BMS-like memories. First, a simplified model where a bald Schwarzschild black hole is considered, and the implantation of supertranslation memory on it is indicated. Next, this simplified model is further improved by considering a rapidly spinning Kerr black hole endowed with a supertranslation hair. In both the cases we analyze the possible way of obtaining imprints of such near-horizon memory in GWs by placing an intermediate mass binary black hole (IMBBH) as proxies for idealized detectors.
II.1 Memory for a non-rotating Black Hole
The detection setup uses a toy model in which an eternal Schwarzschild black hole is endowed with a linearised supertranslation hair by some asymmetric shockwave that fall into it as proposed in Hawking et al. (2017); Strominger (2018). This process implants a linearized supertranslation hair to a bald Schwarzschild black hole:
@@ II.1 Memory for a non-rotating Black Hole
This change in the transverse part of the metric components gives rise to memory for an observer sitting at null infinity. As the physical degrees of freedom are encoded in the transverse part of the metric, this produces a genuine effect.
BMS-like symmetries can also be recovered near the horizon of a black hole that preserves the near-horizon asymptotic structure Donnay et al. (2016b, 2018). To see this, one may expand the metric in Eq. (II.1) near the horizon by a suitably defined radial coordinate $\rho=r-r^{h}_{(f)}$ and relate the parameters of the near-horizon metric with the implanted supertranslation field. To the leading order in $\rho$ this metric reads Donnay et al. (2016b, 2018):
@@
The changes induced via horizon supertranslation (or superrotation) are termed as ‘black hole memory effect’ Donnay et al. (2018). The change induced by the shock-wave can be realized as the action of supertranslation and superroation on the asymptotic metric. From the perspective of a near-horizon observer the changes will get captured via the metric components (II.1). One can obtain the charges corresponding to these near-horizon symmetries. The supertranslation and superrotation charges depend upon the field $f$ or its derivatives, 333see Donnay et al. (2016b) for details. and therefore, if we detect $f$ by any mechanism, we detect the supertranslation or the black hole memory.
@@ II.2 A toy detection model
Figure 1: Schematic representation of inspiralling IMBBH source emitting GWs in the vicinity of the horizon of a bald SMBH (on the left). Once the supertranslation (ST) ‘hair’ is attached to the bald SMBH due to some astrophysical shock wave, the proper distance between the components of the IMBBH is changed. The magnitude of the change in separation (ST memory) denoted by $\Delta L$ is given by Eqs. ( 10,II.3). As the IMBBH separation increases (decreases), the binary will start emitting GWs at a lower (higher) frequency in comparison to the initial configuration (on the left).
@@
It is apparent that to detect the near-horizon supertranslation memory, one needs to place the detectors close to the horizon. However, detection from a far region would not be possible until the detectors can have some mechanism to communicate with the remote region. Now, we propose a scenario (probe for detection) that might capture this supertranslation memory of a Schwarzschild black hole. Suppose that two intermediate-mass black holes (IMBHs) in a binary are in the vicinity of a bald supermassive black hole (SMBH) and they act like detectors. The IMBHs initially continue to generate GWs that are detectable by LISA. Now suddenly the SMBH (almost instantaneously) gets endowed with a (linear) supertranslation hair by sending a shockwave into it as described above. This sudden effect would change the relative proper distance between the inspiraling IMBHs and would ultimately alter the inspiraling frequency and frequency evolution as illustrated in Figure 1. Note that the shockwave acts only as an agent that implants the supertranslation hair on the bald SMBH. This is a convenient way to generate a BH state with a (asymptotic) supertranslation field. However, as already alluded, this changed metric can also be achieved via Lie transporting the bald metric.
@@
We consider the IMBHs to be separated only in the angular directions initially, i.e. they are at the same radial distance from the centre of the SMBH. After the hair is implanted, the separation between the black holes changes. This can be estimated using local coordinates. We will further assume that the supertranslation field is a function of $\theta$ only. This would reduce the algebra and at the same time offer sufficient insight into the desired investigation.
@@
Now, let $L_{0}$ be the initial proper distance (angular) between the test bodies (IMBHs). In that case, the separation would only be governed by the bald angular part of the metric. Therefore $L_{0}^{2}=4M^{2}\gamma_{AB}d\Theta^{A}d\Theta^{B}.$ After the implantation of the supertranslation memory, the changed distance would be governed by the metric given by Eq. (II.1) or Eq.(II.1). If we consider the position of IMBHs almost near the horizon of SMBH, then we should go ahead with the metric in (II.1). However, this situation will practically be untractable from the far region, as the distance between the binary and the horizon of SMBH will be within the ISCO of SMBH. Therefore, we take the metric described in Eq. (II.1) which is valid at any finite distance from the black hole. If $\Delta L$ is the shift in the angular separation, then a straightforward computation yields:
$\Delta L$ $\simeq{r_{0}\partial^{2}_{\theta}f(\delta\theta^{2}-\sin^{2}\theta\delta\phi^{2})\over 2L_{0}}$
$\simeq-{3r_{0}\over 2L_{0}}q\cos(2\theta)(\delta\theta^{2}-\sin^{2}\theta\delta\phi^{2}),$ (9)
@@
where we have chosen $f(\theta)$ to be equal to the second degree Legendre Polynomial $f=qP_{2}(\cos\theta).$ $q$ is a parameter that encapsulates the strength of the supertranslation hair Sarkar et al. (2022); Compère and Long (2016). For the case under consideration, $q$ should be small, as it is part of a weak supertranslation field. For consistency, the initial separation $L_{0}$ should be much greater than the separation $\Delta L$ induced by the field $f$. Here, $\delta\theta$ and $\delta\phi$ are, respectively, the initial polar and azimuthal separations between the objects. The shift in the separation of IMBHs, parametrized by $q$, will induce a change in the frequency of their inspiraling motion. We set IMBBH at distance $r_{0}=(1+0.001)R$ with $R\simeq 3(2M+{1\over 2}D^{2}f)$ as the IMBHs should be at least in the vicinity of ISCO of SMBH. Further, we introduce a dimensionless parameter $q_{0}=q/M$ and rewrite the leading contribution from (II.2) as
$\Delta L\simeq-{9M^{2}\over L_{0}}q_{0}\cos(2\theta)(\delta\theta^{2}-\sin^{2}\theta\delta\phi^{2}).$ (10)
We will use Eq. (10) to demonstrate the detectability of the mis-match in LISA between the original GW signal and the GW signal modulated due to this sudden change in the separation between the IMBHs in the section III.
II.3 A more realistic model
The model considered so far can be improved further. In the previous case, the ISCO of the background Schwarzschild SMBH is not very near to the horizon and one may wonder whether the detection of ST hair can be trusted as a genuine near-horizon memory. In this regard, we now consider a rapidly spinning Kerr black hole as the SMBH. A rapidly spinning Kerr black hole’s metric can be presented as an extreme Kerr metric. Such rapidly spinning SMBHs are indeed expected in nature McClintock et al. (2006); Reynolds (2014); Gou et al. (2014). The extreme Kerr-metric near the horizon is obtained from the traditional Boyer-Lindquist (BL) form by employing several coordinate transformations Bardeen and Horowitz (1999); Chruściel (2015). We start from the BL form:
$ds^{2}=-({d{\bar{r}^{2}}\over\bar{\Delta}}+{d\bar{\theta}^{2}})\bar{\rho}^{2}+(d{\bar{t}}-a\sin^{2}\bar{\theta}d{\bar{\phi}})^{2}{\bar{\Delta}\over\bar{\rho}^{2}}-(\bar{A}^{2}d{\bar{\phi}}-ad{\bar{t}})^{2}{\sin^{2}\bar{\theta}\over\bar{\rho}^{2}},$ (11)
with
$\bar{\Delta}(\bar{r})=\bar{r}^{2}+a^{2}-2M\bar{r},$ (12)
$\bar{\rho}^{2}(\bar{r},\bar{\theta})=\bar{r}^{2}+a^{2}\cos^{2}\bar{\theta},$ (13)
$\bar{A}^{2}(\bar{r})=\bar{r^{2}}+a^{2},$ (14)
where $M,a$ are the mass and angular momentum per unit mass of the black hole. As depicted in Bardeen and Horowitz (1999); Chruściel (2015), we produce a similar kind of metric as in Eq. (II.1) by performing the following transformations first: $\bar{t}=\epsilon^{-1}t^{\prime},\bar{r}=M+\epsilon r^{\prime},\bar{\theta}=\theta^{\prime},\bar{\phi}=\phi^{\prime}+\epsilon^{-1}{a\over r_{0}^{2}}t^{\prime}$, with $r_{0}^{2}=M^{2}+a^{2},$ and taking $\epsilon\to 0$ limit, and then going to a regular advanced null coordinate compatible with the Bondi-Sachs gauge Fletcher and Lun (2003):
$ds^{2}=(1-\sin^{2}\theta)\left[-{R^{2}\over 2M^{2}}dV^{2}+2dVdR+2M^{2}d\theta^{2}\right]$
$+{4M^{4}\sin^{2}\theta\over 2M^{2}-M^{2}\sin^{2}\theta}(d\Phi+{R\over 2M^{2}}dV)^{2}\,$, (15)
where $\Phi=\phi^{\prime}-\log(R/\sqrt{2}M),$ and $t^{\prime}=V-2M/R$. Here we have set $M=|a|$ for the extreme case. This metric also played an important role in exploring the emergence of conformal symmetries near the horizon of a highly spinning black hole Guica et al. (2009); Porfyriadis and Strominger (2014). After supertranslating this metric, we get Galoppo et al. (2024):
$ds^{2}=g_{VV}dV^{2}+g_{VR}dVdR+g_{V\Phi}dVd\Phi+g_{V\theta}dVd\theta$
$+g_{\theta\theta}d\theta^{2}+g_{\Phi\Phi}d\Phi^{2}\,$, (16)
where,
$g_{\theta\theta}=(2M^{2}-M^{2}\sin^{2}\theta)\left(1-{2\over M}\partial^{2}_{\theta}f\right)+M\cos 2\theta\partial_{\theta}f\,$,
and
$g_{\Phi\Phi}=\left[{4M\over(2-\sin^{2}\theta)}\right]$ $\times$
$\left[M\sin^{2}\theta-{2\sin 2\theta\over 1-\sin^{2}\theta}\partial_{\theta}f-\partial^{2}_{\Phi}f\right]\,.$
We do not need the explicit forms of the other metric components for the computation of the memory. The horizon of the supertranslated extreme-Kerr is at $R_{h}=M+{1\over 2}D^{2}f.$ As we know, a class of ISCOs (for prograde orbits) for an extreme Kerr black hole is situated almost at the horizon Bardeen et al. (1972). Again, computing the shift in angular separation of the binary components situated close to the ISCO (say $r_{0}=(1+0.001)R$), with ISCO at $R\sim M+\mathcal{O}(f)$, we get the following:
$\Delta L\simeq-{1\over 4L_{0}}M^{2}q_{0}(12\cos(2\theta)+3\cos(4\theta)+1)\cos(2\phi)\delta\theta^{2}$
$+{32M^{2}q_{0}\sin^{2}(\theta)\cos(\theta)\cos(2\phi)(5\sin(\theta)+\sin(3\theta)-2\cos(\theta))\over(\cos(2\theta)+3)^{2}L_{0}}\delta\phi^{2}.$ (17)
444we have used $\phi$ instead of $\Phi$ for this expression.
Here, we have considered $f=qf(\theta,\phi)=q{\rm Re}(Y_{22})$, where ${\rm Re}(Y_{22})$ is the real part of $l=m=2$ spherical harmonic $Y_{lm}(\theta,\phi)$.
III Frequency modulation of the binary